OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (anonymous):

these are my choices y = 3x + 4 y = one-thirdx + 3 y = –3x – 6 y = negative one-thirdx – 5 @CarlosGP

OpenStudy (anonymous):

Try it yourself with instructions given: for Y=mX+B, the perpendicular has to be: Y=(-1/m)X+C, calculate the slope -1/m and the value C(substituting the values of the point given)